PATHALIAS(8)  UNIX Programmer’s Manual  PATHALIAS(8)

”Public Domain” NAME

pathalias, makedb, arpatxt – mail routing tools SYNOPSIS

pathalias [ –ivcDf ] [ –t link ] [ –l host ] [ –d link ] [ –g file ] [ files ... ] makedb [ –a ] [ –o dbmfile ] [ files ... ] arpatxt [ –@fi ] [ –g gateway ] [ –p privatefile ] [ –d directory ] [ files ... ]

DESCRIPTION Pathalias computes the shortest paths and corresponding routes from one host (computer system) to all other known, reachable hosts. Pathalias reads host-to-host connectivity information on standard input or in the named files, and writes a list of host-route pairs on the standard output. Here are the pathalias options: –i Ignore case: map all host names to lower case. By default, case is significant. –c Print costs: print the path cost before each host-route pair. –v Verbose: report some statistics on the standard error output. –D Terminal domains: see domains section. –f First hop cost: the printed cost is the cost to the first relay in a path, instead of the cost of

the path itself; implies (and overrides) the –c option. –g file Dump the edges of the graph into the named file. –l host Set local host name to host. By default, pathalias discovers the local host name in a system-

dependent way. –d arg Declare a dead link, host, or network. If arg is of the form “host-1!host-2,” the link from

host-1 to host-2 is treated as an extremely high cost (i.e., DEAD link. If arg is a single host name, that host is treated as dead and is used as a relay host of last resort on any path. If arg is a network name, the network requires a gateway.

–t arg Trace input for link, host or network on the standard error output. The form of arg is as

above. Makedb takes pathalias output and creates or appends to a dbm(3) database. Here are the makedb options: –a Append to an existing database; by default, makedb truncates the database. –o dbmfile Identify the output file base name. Arpatxt converts the Internet hosts table hosts.txt into pathalias input. Here are the arpatxt options: –@ Generate pathalias input that specifies ‘@’ style addressing. The default is to produce

pathalias input that specifies ‘!’ style addressing. –f “Filter mode” — write output on stdout. Normally, arpatxt writes the description of each

domain into a separate file. –i Map output to lower case.

–g arg Declare a gateway to the Internet or one of its subdomains. If arg contains one or more dots, the left-hand side component that contains no dots is declared a gateway to the domain to the right of the dot. Otherwise, arg is declared a gateway to the Internet as a whole.

–p privatefile Privatefile contains a list of host names that conflict with other names.

–d directory Write output files in directory. A line beginning with white space continues the preceding line. Anything following ‘#’ on an input line is ignored.

A list of host-to-host connections consists of a “from” host in column 1, followed by white space, followed by a comma-separated list of “to’ hosts, called links. A link may be preceded or followed by a network character to use in the route. Valid network characters are ‘!’ (default), ‘@’, ‘:’, and ‘%’. A link (and network character, if present) may be followed by a “cost” enclosed in parentheses. Costs may be arbitrary arithmetic expressions involving numbers, parentheses, ‘+’, ‘–’, ‘*’, and ‘/’. Negative costs are prohibited. The following symbolic costs are recognized:

LOCAL25 (local-area network connection) DEDICATED95 (high speed dedicated link) DIRECT200 (toll-free call) DEMAND300 (long-distance call) HOURLY500 (hourly poll) EVENING1800 (time restricted call) DAILY5000 (daily poll, also called POLLED WEEKLY30000 (irregular poll)

In addition, is a very large number (effectively infinite), and are –5 and +5 respectively, for baud-rate or quality bonuses/penalties, and is –80, for adjusting costs of links that use high-speed (9.6 Kbaud or more) modems. These symbolic costs represent an imperfect measure of bandwidth, monetary cost, and frequency of connections. For most mail traffic, it is important to minimize the number of hosts in a route, thus, e.g., * 24 is much larger than If no cost is given, a default of 4000 is used. For the most part, arithmetic expressions that mix symbolic constants other than and make no sense. E.g., if a host calls a local neighbor whenever there is work, and additionally polls every evening, the cost is not Some examples:

down princeton!(DEDICATED, tilt, %thrash(LOCAL princeton topaz!(DEMANDLOW topaz @rutgers(LOCAL1)

If a link is encountered more than once, the least-cost occurrence dictates the cost and network character. Links are treated as bidirectional but asymmetric: for each link declared in the input, a reverse link is assumed. If the “to” host in a link is surrounded by angle brackets, the link is considered terminal, and further links beyond this one are heavily penalized. E.g., with input

seismo <research>(10), research(100), ihnp4(10) research allegra(10) ihnp4 allegra(50)

the path from seismo to research is direct, but the path from seismo to allegra uses ihnp4 as a relay, not research. The set of names by which a host is known to its neighbors is called its aliases. Aliases are declared as follows: name = alias, alias ...

The name used in the route to or through aliased hosts is the name by which the host is known to its predecessor in the route. Fully connected networks, such as the or a local-area network, are declared as follows: net = {host, host, ...} The host-list may be preceded or followed by a routing character (‘!’ default), and may be followed by a cost (default 4000). The network name is optional; if not given, pathalias makes one up.

etherhosts = {rahway, milan, joliet}!(LOCAL ringhosts = @{gimli, alida, almo}(DEDICATED = {etherhosts, ringhosts}(0)

The routing character used in a route to a network member is the one encountered when “entering” the network. See also the sections on gateways and domains . Connection data may be given while hiding host names by declaring private {host, host, ...} Pathalias will not generate routes for private hosts, but may produce routes through them. The scope of a private declaration extends from the declaration to the end of the input file in which it appears, or to a private declaration with an empty host list, whichever comes first. The latter scope rule offers a way to retain the semantics of private declarations when reading from the standard input. Dead hosts, links, or networks may be presented in the input stream by declaring dead {arg, ...} where arg has the same form as the argument to the –d option. To force a specific cost for a link, delete all prior declarations with delete {host-1!host-2} and declare the link as desired. To delete a host and all its links, use delete {host} Error diagnostics refer to the file in which the error was found. To alter the file name, use file {filename} Fine-tuning is possible by adjusting the weights of all links from a given host, as in adjust {host-1, host-2(LOW), host-3(–1)} If no cost is given a default of 4000 is used. Input from compressed (and uncompressed) files can be piped into pathalias with the following script.

for i in $*; do case $i in *.Z) echo ”file {‘expr $i : ’.Z’‘} zcat $i ;; *) echo ”file {$i}” cat $i ;; esac echo ”private {}” done

A list of host-route pairs is written to the standard output, where route is a string appropriate for use with printf(3), e.g.,

rutgers princeton!topaz!%s@rutgers

The “%s” in the route string should be replaced by the user name at the destination host. (This task is normally performed by a mailer.) Except for domains, the name of a network is never used in routes. Thus, in the earlier example, the path from down to up would be “up!%s,” not “princeton-ethernet!up!%s.” A network is represented by a pseudo-host and a set of network members. Links from the members to the network have the weight given in the input, while the cost from the network to the members is zero. If a network is declared dead, the member-to-network links are marked dead, which effectively prohibits access to the network from its members. However, if the input also shows an explicit link from any host to the network, then that host can be used as a gateway. (In particular, the gateway need not be a network member.) E.g., if is declared dead and the input contains

CSNET= {...} csnet-relay CSNET then routes to hosts will use csnet-relay as a gateway. A network whose name begins with ‘.’ is called a domain. Domains are presumed to require gateways, i.e., they are DEAD The route given by a path through a domain is similar to that for a network, but here the domain name is tacked onto the end of the next host. Subdomains are permitted. E.g.,

harvard .EDU# harvard is gateway to .EDU domain .EDU= {.BERKELEY .UMICH .BERKELEY= {ernie}

yields

ernie ...!harvard!ernie.BERKELEYEDU%s

Output is given for the nearest gateway to a domain, e.g., the example above gives

.EDU...!harvard!%s

Output is given for a subdomain if it has a different route than its parent domain, or if all its ancestor domains are private. If the –D option is given on the command line, pathalias treats a link from a domain to a host member of that domain as terminal. This property extends to host members of subdomains, etc, and discourages routes that use any domain member as a relay. Makedb builds a dbm(3) database from the standard input or from the named files. Input is expected to be sequence of records, each consisting of a key field and a data field separated by a single tab. If the tab is missing, the data field is assumed to be empty.

FILES ET AL.

/usr/local/lib/palias.{dir,pag} default dbm output newsgroup comp.mail.maps likely location of some input files getopt(3), available from comp.sources.unix archives (if not in the C library).

BUGS

The –i option should be the default. The order of arguments is significant. In particular, –i and –t should appear early. Pathalias can generate hybrid (i.e. ambiguous) routes, which are abhorrent and most certainly should not be given as examples in the manual entry. Experienced mappers largely shun ‘@’ when preparing input; this is historical, but also reflects UUCPs facile syntax for source routes. Multiple ‘@’s in routes are loathsome, so pathalias resorts to the “magic %” rule when necessary. This convention is not documented anywhere, including here. The –D option elides insignificant routes to domain members. This is benign, perhaps even beneficial, but confusing, since the behavior is undocumented and somewhat unpredictable.

SEE ALSO P. Honeyman and S.M. Bellovin, “PATHALIASor The Care and Feeding of Relative Addresses,” in Proc. Summer USENIXConf., Atlanta, 1986.